OpenStudy (anonymous):

what is a line perpendicular to Fx=[(-e^-1)*x-(e^-1)]?

zepdrix (zepdrix):

\[\Large\rm F(x)=-x e^{-1}-e^{-1}\]Am I reading that correctly? Something like this?

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

I'll put the x back behind like you had it written. \[\Large\rm F(x)=\color{orangered}{\left(-e^{-1}\right)}x-e^{-1}\]It will make more sense in this form. Remember your slope-intercept form of a line?\[\Large\rm y(x)=\color{orangered}{m}x+b\]

zepdrix (zepdrix):

The coefficient on our x is the slope, we can recognize that from the formula, yes?

OpenStudy (anonymous):

i think so but im not sure since its part of a longer question

zepdrix (zepdrix):

For a line to be `perpendicular` to this one, it means the slope will be the `negative reciprocal` of this slope. So we have some other line that we're trying to form:\[\Large\rm y=mx+b\]And the slope will be the negative, and flip, of our slope from F(x).

OpenStudy (anonymous):

okay, i get that but i get confused when working with e

zepdrix (zepdrix):

Yes this is going to be weird :) We're working with `e`, and we have to pay attention to our exponent rules very carefully in this next step.

zepdrix (zepdrix):

So we take our slope, \[\Large\rm -e^{-1}~negative\to\qquad =e^{-1}~and~flip\to\qquad=\frac{1}{e^{-1}}\]

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

Then we want to remember our exponent rule, to take the negative off of the power we toss him up into the numerator, yes?\[\Large\rm \frac{1}{e^{-1}}=\frac{e^{+1}}{1}\]

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

Ignore the denominator, since it's just one. Ignore the exponent since it's just one... So our new slope, our perpendicular line's slope, is simply \(\Large\rm m=e\)

zepdrix (zepdrix):

Fancy stuff huh? :O

OpenStudy (anonymous):

oh thanks XD

zepdrix (zepdrix):

c: